ANALYSIS OF THE VIBRATIONAL SPECTRA OF P_{3}N_{3}(OCH_{2}CF_{3})_{6} AND P_{4}N_{4}(OCH_{2}CF_{3})_{8}

\underline{ADRIAN K. KING}, DAVID F. PLANT, PETER GOLDING, Atomic Weapons Establishment, Aldermaston, Berkshire, RG7 4PR, United Kingdom; MICHAEL A. LAWSON and PAUL B. DAVIES, University of Cambridge, Department of Chemistry, Lensfield Road, Cambridge, CB2 1EW, United Kingdom.

The cyclic phosphazene trimer P_{3}N_{3}(OCH_{2}CF_{3})_{6} and the related cyclic tetramer P_{4}N_{4}(OCH_{2}CF_{3})_{8} have been proposed as the major low-to-medium temperature pyrolysis products of the parent polyphosphazene (PN(OCH_{2}CF_{3})_{2})_{n} ^{a}^{b}. Recently, both molecules have been synthesized, isolated and their vapour-phase vibrational spectra recorded using a high-resolution FTIR instrument.

The interpretation of these spectra is achieved primarily by comparison with the results of high-quality density functional calculations, which enable the principal absorption features to be assigned and conclusions to be drawn regarding the geometries and conformations adopted by both molecules. These in turn allow interesting comparisons to be made with analogous cyclic halo-phosphazenes such as P_{3}N_{3}Cl_{6} and P_{4}N_{4}Cl_{8} ^{c}^{d}.

Work to record in situ the spectra of the vapour-phase pyrolysis products of (PN(OCH_{2}CF_{3})_{2})_{n} and to analyse these results in terms of the tetramer and trimer spectra will also be presented.

aS. V. Peddada and J. H. Magill Macromolecules 16 (1983) 1258-1264.

bH. R. Allcock, G. S. McDonnell, G. H. Riding, and I. Manners Chem. Mater. 2 (1990) 425-432.

cT. R. Manley and D. A. Williams Spectrochimica Acta 23A (1966) 149-165.

dV. Varma, J. R. Fernandez and C. N. R. Rao J. Mol. Struct. 198 (1989) 403-412.